ars with an accumulated portfolio valus that must be invested at the end of each an annual rate of return of 6% 3. A man planplease explain
Answer #1

In case of any issue please ask in the comments

3.
a.

In this case Years or n = 35, interest or i = 6%, Future value = 1,500,000

Using the Annuity formula

PMT = 13,460.8

b.

=PMT(6%,35,,1500000)

Answer would be 13,460.79 per annum

4.

a.

PV = FV X 1 / (1 + r)^n = 80,000 X 1/(1+4.75%/12)^144 = 80,000 X 0.566162 = $ 45,292.97

b.

=PV(4.75%,12,,80000)

Answer would be $ 45,839.68